OpenStudy (anonymous):

Use substitution to solve the following system of equations. Enter your answer as an ordered pair, like this example: (4, -5). y=x-19 2x-y=27

OpenStudy (anonymous):

y = x - 19 <-------- use this in the other equation. 2x - y = 27 2x - (x - 19) = 27 2x - x + 19 = 27 x + 19 = 27 x = 8 Don't forget to find y. y = x - 19 = (8) - 19 = -11 So your answer is (8, -11)